Problem 2. Find the general solution of the nonhomogeneous advancement operator equation (A - 5)(A + 2)f(n) = 3
Answer #1

lo (4.5) (A+2) f(n) = 3h Characterstic ean (7-5) (7+2) = 0 2=5, -2 So Homogenous solh. f(n) = (5) + C (-2) Particular solutio